Subject: 3.5-m Users Committee minutes 5/01/06
From: Bruce Gillespie
Submitted: Fri, 5 May 2006 10:20:17 -0600
Message number: 963
(previous: 962,
next: 964
up: Index)
APO 3.5-m Users Committee Phonecon, 5/01/06
Attending: Suzanne Hawley, Bruce Gillespie, Jon Holtzman, Remy
Indebetouw, Michael Strauss, John Wilson, John Bally, Bruce Balick,
Don York, Karl Glazebrook
Absent: Al Harper, Russet McMillan
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Princeton (Michael Strauss): The Princeton users are happy. As to
the Target-of-Opportunity questions posed at the last meeting,
Michael said there is strong PU support for the idea of promoting ToO
use of the telescope--it's what the 3.5-m telescope excels at, plays
to our strength, but that ToO programs should be vetted and held to
the same science standards as regular observing programs. Suzanne
added that unplanned last-minute ToO proposals will be considered
without a written proposal, but in these cases she expects one to be
submitted rapidly (within 1-2 days) if the ToO program will be
ongoing. Such proposals need to submitted through the institutional
scheduler as discussed at the last users committee meeting.
Colorado (John Bally): No CU user feedback to report, and John also
feels strongly that ToO programs should be supported.
NMSU (John Holtzman): Nothing to report from users or on the ToO
questions.
Johns Hopkins (Karl Glazebrook): Only complaint is about the weather
[editor's comment: JHU observers must have been unlucky recently],
and Karl said he got no user response to the ToO issue.
Chicago: (Don York for Al Harper): Don said that the UC faculty
strongly supports ToO use of the telescope, and he also agrees that a
proposal and review should be required for ToO programs.
UWashington (Bruce Balick): Nothing to report from the UW users.
Univ. of Virginia (Remy Indebetouw): UVa users are generally happy,
but Remy has received no user feedback to the ToO questions.
**********************************
Discussion of telescope/instruments report:
BruceG summarized highlights from the report (see below). BruceG
also mentioned that there was an accident on the APO road last week
where two electric company linemen were killed in a fall while
attempting to replace a defective power pole. Several staff members
from APO and NSO/Sunspot were first responders at the accident scene,
and the deaths of two young men, both from Cloudcroft, is a terrible
tragedy for the community.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights,
3/30/06 through 4/27/06
Bruce Gillespie
0) Overview
Highlights include continued dry conditions at the site (good for
observing, bad for forest-fire risk), a well-behaved telescope and
instruments, good progress on the new top-end and DIS-upgrade projects.
1) Telescope
The 3.5-m telescope was again largely well behaved during the past
month and on the sky nearly every night. An investigation into why
we occasionally see slightly elongated images has suggested a problem
with one of the tertiary mirror actuators. Replacing a suspect
encoder did not solve the problem, and we believe that the tertiary
will need to be removed to find the root cause, which is probably a
damaged flex pivot. In the meantime, we have turned off the servo
that does fine alignment of the tertiary, in an attempt to eliminate
residual image wander at the expense of slightly degrading pointing
accuracy. Also, a new version of the TCC software was installed and
tested, a pointing model was made, and a method to improve the speed
of computed telescope offsets was successfully tested. Lastly, there
was in unexplained sudden occurrence of a telescope pointing glitch
of a few arcminutes--this was corrected for in a subsequent pointing
model, but the root cause remains a mystery.
2) Instruments
The performance of all the instruments was generally good.
Noteworthy was the installation and successful testing of a new
control computer for SPIcam.
3) Engineering and CIF projects
Design and fabrication work continued on the new top end, and the
project remains on track. The DIS upgrade project (gratings, optics,
new detector) continued; parts are being ordered and components are
being machined, and a schedule for their installation was drafted.
The new red detector has still not been delivered from e2v.
4) Miscellaneous
The dryness, wind, and onset of the pine-pollen season has elevated
the dust/particulate levels at APO to nearly double the normal
levels. We are liquid-CO2 cleaning the telescope optics several
times a week (which has resulted in LC02 shortages from our vendor)
in an attempt to keep the telescope on the sky as much as practical.
We have completed the fire-protection enhancements to the site,
including outside water sprinklers, tree removals, a fire truck on
station, leak repairs to our water-storage tank, and the installation
of exterior metal shutters on building windows.
**********************************
DIS upgrade plans - Hawley, Holtzman:
JonH reported that it's time that the users became aware of the soon-
to-happen upgrades to DIS. In mid-June, two new grating and mounting
assemblies will be installed. On both the red and blue sides of the
instrument, the medium and low gratings are being replaced with new
medium gratings that we hope become the "standard" for low/medium
resolution. The blue-high grating is not being changed, but a new
red-high grating is being added sometime later to replace/supplement
the existing red-high. All the old gratings can be swapped back into
DIS on request, but this should be done sparingly, if at all, because
the new gratings should give better overall performance if they work
as advertised.
Jon said he'd post a fact sheet about the new gratings shortly so
that users can learn what the changes will be. Later in June, we
plan to install a new prism into the blue side of the spectrograph to
recover UV sensitivity (hopefully to the atmospheric cutoff), and
replace the red detector with a thick-substrate version that should
reduce or eliminate the red-fringing problem.
**********************************
Visiting instrument proposal - Hawley:
Suzanne reported that we have received a proposal from a non-ARC
astronomer to build a wide-field NIR camera and grism spectrometer
with a 15-arcminute FOV and a programmable slit mask, and bring the
instrument to the 3.5-m for their and our use. This would be in
exchange for telescope time much like the arrangements we currently
have for CorMASS and the Goddard Fabry-Perot instruments. Several of
the Users Committee members expressed strong support for continuing
to pursue this initiative, and commented on how such an instrument
would complement NIC-FPS and further our interests in general NIR
science capabilities.
ACTION: Suzanne asked each Users Committee member to send her a
paragraph or two in the next two days with an expression of interest
and/or questions regarding this proposed NIR instrument.
**********************************
Futures Committee news - York:
Don will send BruceG copies of two Futures Committee documents that
will be forwarded to the Users Committee members, who are then to
help their institutional Futures Committee member share them with
their constituent faculties and collect feedback. Don reported that
the Futures Committee report will recommend that the 3.5-m telescope
continue to be operated after 2008 much the same as now, with new
instruments, remote observing, CIF funding at current levels, and
with no general repurposing needed to remain viable and productive.
For the 2.5-m telescope, there is now a substantial list of post-2008
proposals, which Don summarized, to use the wide-field capabilities
of the telescope to do a host of interesting and largely unique
science projects, some with the existing instrumentation and others
requiring the advent of new instruments. In figuring out which of
these programs to push forward, there are difficulties relating to
the continuing lateness of the NSF Senior Review, and in blending
several of these proposals into an integrated program like SDSS-II;
it will take months and lots of back and forth dialog within ARC to
resolve this. Furthermore, the committee will recommend
consideration of restructuring ARC in a manner that can allow
continued operation of both telescopes in the absence of agency
funding support.
ACTION: BruceG to forward Futures documents to Users Committee
members, who then help the institutional Futures Committee member
share them with their faculties and gather feedback.
**********************************
Previous Meetings Action Items report:
[open from previous months]:
ACTION: The Users Committee members will need to help disseminate
the [Futures Committee] draft report to their users. STATUS: Draft
coming soon.
ACTION: BruceG will ask Russell Owen what the pros and cons are of
keeping TUI compatible with older versions of operating systems and
libraries. STATUS: Discussed with RO, no final decision yet.
ACTION: Users Committee members should poll their users to ask if
there are compelling reasons to keep TUI operable without using the
latest version of operating systems. STATUS: Open, only little
feedback so far.
ACTION: [re Triplespec white paper] The committee members are to
forward any feedback to Suzanne within a week, from which she will
report our assessment to the ARC Board of Governors. STATUS:
Completed.
[new actions from last month]:
ACTION: BruceG to discuss issues of computer and network security
with Fritz Stauffer. STATUS: Working group (staff and invited
users) being formed to frame requirements.
ACTION: UC members should poll their constituents about ToO
observations on the telescope. Do they envision doing this type of
science? Would expanding ToO opportunities be useful to them? How do
they feel about being asked to voluntarily give up time, and to
having their time pre-empted (perhaps with later payback)? STATUS:
Discussed in 5/1 Users Committee meeting.
ACTION: BruceG will ensure that the orientation checklist is up to
date and available on-line, so that prospective new users can review
it prior to their orientation site visit. STATUS: Open, plan to
complete in early May.
ACTION: Jon Holtzman to talk to Fritz Stauffer and BruceG about
installing a RAID backup at NMSU. STATUS: Discussed, system
components ordered.
**********************************
Next Users Committee phonecon will be on Monday, 12 June, at 8:30 AM
Pacific Time. Agenda and other materials will be sent to the
committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 963 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO